Marlene: Using the following schedule, define the equilibrium price and quantity. Describe the situation at a price of $7. What will occur? Describe the situation at a price of $3. What will occur?

Price ($)

Quantity Demanded

Quantity Supplied

1

1000

120

2   

985

200

3

850

315

4

720

400

5

600

450

6   

520

520

7   

320

650

8   

200

718

9

180

825

10   

100

997
Answer #1

Equilibrium price will be the price at which quantity demanded equals quantity supplied.
From the table we can see that when,
Price = $6 => Quantity demanded = Quantity Supplied = 520
Therefore, at equilibrium P=$6 Q=520

At P=$7, quantity demanded(Qd) < quantity supplied (Qs).
Therefore, there is excess supply of Qs - Qd = 650 - 320 = 330 units in the market. Therefore, price has to decrease so that Qd increases and Qs decreases and the market reaches at equilibrium.

At P=$3, quantity demanded(Qd) > quantity supplied (Qs).
Therefore, there is excess demand of  Qd - Qs  = 850 - 315 = 535 units in the market. Therefore, price has to increase so that Qd decreases and Qs increases and the market reaches at equilibrium.